A grocer mixed peanuts(whose price is 90 cents per kilogram)with almonds(whose price is $1.10 per kilogram)to make a 20-kilogram mixture of nuts for sale at 98 cents per kilogram. How many kilograms of peanuts and how many kilograms of almonds are in the mixture?: A grocer mixed peanuts(whose price is 90 cents per kilogram)with almonds(whose price is $1.10 per kilogram)to make a 20-kilogram mixture of nuts for sale at 98 cents per kilogram. How many kilograms of peanuts and how many kilograms of almonds are in the mixture?
Answer by mangopeeler07(429)

With two variables------x=number of kilograms of 90cent peanuts and y=number of kilograms of 110cent almonds.
Set up an equation using the word problem: 90x+110y=98(20) because the coefficients are the prices of the almonds and peanuts and the variable and 20 are the number of them. Then set up another equation x+y=20 because the sum of the kilograms is 20. Then solve for x and get x=20-y. Plug that in the first equation and get 90(20-y)+110y=98(20). Then distribute and get 1800-90y+110y=1960. Combine like terms and get 1800+20y=1960. Subtract 1800 from both sides and get 20y=160. Divide both sides by 20 and get y=8. Then plug 8 in for y in the second equation and get x+8=20. So x=12. So 8 kilograms of 110cent almonds and 12 kilograms of 90 cent peanuts is your final answer.


With one variable------x=number of kilograms of 90cent peanuts and 20-x=number of kilograms of 110cent almonds.90x+110(20-x)=98(20)
Distribute and get 90x+2200-110x=1960. Then combine like terms and get 2200-20x=1960. Subtract 2200 from each side and get -20x=-240. Divide both sides by 20 and get x=12. This is the number of kilograms of 90 cents peanuts. Then do 20-x or 20-12 and get 8. This is the number of kilograms of 110cent almonds. So 8 kilograms of 110cent almonds and 12 kilograms of 90 cent peanuts is your final answer.
Question 147690: A grocer mixed peanuts(whose price is 90 cents per kilogram)with almonds(whose price is $1.10 per kilogram)to make a 20-kilogram mixture of nuts for sale at 98 cents per kilogram. How many kilograms of peanuts and how many kilograms of almonds are in the mixture?: A grocer mixed peanuts(whose price is 90 cents per kilogram)with almonds(whose price is $1.10 per kilogram)to make a 20-kilogram mixture of nuts for sale at 98 cents per kilogram. How many kilograms of peanuts and how many kilograms of almonds are in the mixture?
Answer by stanbon(18060)

A grocer mixed peanuts(whose price is 90 cents per kilogram)with almonds(whose price is $1.10 per kilogram)to make a 20-kilogram mixture of nuts for sale at 98 cents per kilogram. How many kilograms of peanuts and how many kilograms of almonds are in the mixture?
--------------
EQUATIONS:
Quantity: p + a = 20
Value: 90p + 110a = 98*20
-------------------------
Rearrange:
9p + 11a = 196
9p + 9a = 180
------------------
2a = 16
a = 8 (# of pounds of almonds in the mixture)
--------------
p + 8 = 20
p = 12 (# of pounds of peanuts in the mixture)
================
Cheers,
Stan H.